OpenStudy (anonymous):

Find the equation of the line that passes through the point whose coordinates are (1,-1) and has slope 2.

OpenStudy (anonymous):

\[y-y_a=m(x-x_a)\] \[ m=slope \]\[ (x_a,y_a)=point \]

OpenStudy (anonymous):

you are given the point and slope so plug that into the first equation that I wrote that and it the equation of the line. Any questions?

OpenStudy (anonymous):

example: point is (0,1) and slope = -2, so: y - 1 = (-2)(x-0). Then, y - 1 = -2x. Isolating "y" on the left side, y = -2x+1.